# TriviaHG: A Dataset for Automatic Hint Generation from Factoid Questions

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: f the 47th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 14–18, 2024; Washington D.C., USAisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Retrieval models and rankingccs: Information systems Evaluation of retrieval results
@@ 1. Introduction
Automatic question answering (QA) systems (Abdel-Nabi et al., 2023; Rogers et al., 2023) have recently gained significant attention, allowing users to pose questions and receive direct answers. Traditional QA systems excel in specific knowledge domains by extracting relevant text spans from passages. In contrast, Large Language Models (LLMs) (Team et al., 2023; Workshop et al., 2022; Brown et al., 2020a), pre-trained on vast datasets, offer a more versatile approach. LLMs can handle multiple NLP tasks, including automatic QA, and can answer questions across diverse domains. Unlike extractive QA systems, they generate answers by creating new content, showcasing high proficiency and potential as an alternative to traditional QA systems.
Figure 1. Example hints for two sample questions in TriviaHG with their computed convergence (CON) and familiarity (FAM) attributes provided on a scale of 1 to 5 (1 is the lowest and 5 is the highest quality.
@@
The use of hints to assist users is not a new concept in computer science. For example, in the field of automatic hint generation for programming (AHGP)  (McBroom et al., 2021), the task is to generate hints automatically to aid students in coding exercises. However, there is limited research on automatic hint generation specifically for factoid questions (AHGQ) - questions that can be answered with facts expressed in a few words (Jurafsky, 2000; Jatowt et al., 2023). While both AHGP and AHGQ aim to assist humans, the former is more tailored to coding exercises, often suggesting code edits, and the latter is a more general task applicable across various types of questions. Furthermore, there is an additional benefit from focusing on AHGQ task. Hints for questions could be used to enhance Retrieval Augmented Generation (RAG) solutions (Gao et al., 2023; Abdallah and Jatowt, 2023) and Query Expansion (Wang et al., 2023) methods. Specifically, in the context of Retrieval Augmented Generation, incorporating hints as query explanations has the potential to enhance the model’s ability to understand and process queries. As mentioned above, the hints may also find applications in entertainment, such as in game shows like Jeopardy.
As a novel task, AHGQ faces challenges such as the absence of datasets and reliable evaluation methods. Addressing this gap, we introduce a framework for AHGQ and utilize it to generate TriviaHG [^1] The code, dataset, and experimental results are freely available at https://github.com/DataScienceUIBK/TriviaHG, a dataset for automatic hint generation for factoid questions. Our approach relies on web pages as the primary information source and utilizes Bing Chat AI [^2] https://chat.bing.com to generate hints based on online sources, ensuring accuracy and relevance. The hints are designed to be user-friendly and helpful.
As a second contribution, we propose an automatic evaluation method to assess hint quality, considering aspects of convergence and familiarity. The convergence quality is a characteristic that reflects the extent to which a hint can narrow down or eliminate potential answers to a given question. The familiarity is a characteristic that signifies the degree to which the entities mentioned in a hint are well-known or widely recognized. Figure 1 illustrates two questions along with a few examples hints from TriviaHG along with the computed convergence and familiarity quality attributes.
@@ 2.1. QA Task
The previous methods do not focus on generating hints but rather on providing the final answers.
@@ 2.3. Evaluation Methods
Evaluation methods for AHGP can be broadly classified into four categories: (1) User Studies and Surveys, (2) Comparisons with Experts, (3) Analysis Using Historical Data, and (4) Technical Evaluations (McBroom et al., 2021). The User Studies and Surveys category involves collecting feedback from users (typically students) who interact with the hint system (Head et al., 2017; Lee et al., 2018). The Comparisons with Experts category involves comparing the hints generated by the system with those provided by human experts (Piech et al., 2015; Price et al., 2017). The Analysis Using Historical Data involves using past student interaction data to evaluate the effectiveness of the hints (Lazar et al., 2017; Haldeman et al., 2018). The Technical Evaluations category involves assessing the system’s performance in terms of computational efficiency, scalability, and robustness (Marin et al., 2017).
All the evaluation methods for QA are automatic, but there are a few automatic evaluation methods for AHGP. It shows that assessing hint quality automatically is a complex task. However, the automatic evaluation methods for AGHP cannot be employed to assess AGHQ because they just focus on the programming aspects of hints. As said, Jatowt et al. (2023) is the only study focused on the AGHQ, but they also evaluate the quality of hints manually by asking humans to assess the hints and do not propose any automatic evaluation methods for AGHQ. In our work, we propose a fully automated method for evaluating hint qualities.
@@ 3.1.1. Question Selection
We select the TriviaQA dataset (Joshi et al., 2017) as our primary question source because it offers challenging trivia-style questions covering a wide range of diverse topics. It has also been frequently used in QA research (Beltagy et al., 2020; Zaheer et al., 2020; Su et al., 2024; Prasad et al., 2023; Nassiri and Akhloufi, 2023). We filter out questions with fewer than 6 words or exceeding 20 words, as well as those lacking question marks. Additionally, questions whose answers do not have corresponding Wikipedia pages are excluded, as Wikipedia is utilized for evaluating familiarity quality. After applying these criteria, 66,300 questions remain available for further processing.
@@ 3.2.1. Hint Generator
We utilize LLMs to generate hints for the questions, specifically selecting Bing Chat AI as our preferred choice. Bing Chat AI functions as a RAG system, leveraging also information from retrieved documents on the Internet to formulate answers to user questions. This system not only generates answers based on existing online documents but also provides the sources of the hints and answers. In contrast, LLMs like GPT-4 (Achiam et al., 2023), Gemini (Team et al., 2023), and WizardLM (Xu et al., 2023) rely on their stored knowledge to generate answers without indicating the sources. Therefore, Bing Chat AI is considered more reliable for hint generation as it transparently presents the sources of the generated hints. To generate hints for questions, we follow these steps:
@@
    We first prompt Bing Chat AI to answer a question. This enables us to determine if Bing Chat AI is capable of retrieving the answer. If it is unable to do so, we discard the question.
@@
    We instruct Bing Chat AI to produce 10 hints for the question without including the answer in the hints. This approach allows us to generate hints without the risk of answer leakage, i.e., the case when a hint explicitly reveals the answer to the question in its content [^6] For instance, for the question Who won the FIVB Volleyball Men’s World Championship in 2018?, the hint Their official language is Polish. exhibits answer leakage.. Figure 3 illustrates the trend of using prompts in Bing Chat AI for hint generation.
Figure 3. The hint generation system initiates by prompting the question to Bing ChatAI. Then, it produces a snippet serving as the answer. Following this, we assess the correctness of the provided answer. If the answer is correct, we prompt Bing Chat AI to generate 10 hints. The numbers in the brackets denote the source pages for both the answer and the hints.
@@ 3.2.2. Hint Filtering
While we advise Bing Chat AI to exercise caution regarding the problem of answer leakage, there are instances where it might produce hints inadvertently revealing the answer. Additionally, there are occasions when it generates hints that merely rephrase or describe the questions. For instance, for the question In which city are the headquarters of the International Monetary Fund?, the model may generate a hint like The headquarters of the International Monetary Fund is located in this city. Unfortunately, these hints merely repeat the question and do not help users in finding the answer. To address these issues, we follow these steps:
@@
Convergence Evaluation value of convergence
Familiarity Evaluation value of familiarity
3.2.3. Hint Labeling
@@ 4. Automatic Evaluation Method
In this work, we focus on automating the evaluation of convergence and familiarity quality attributes, as, in our opinion, they are the strongest indicators of hint quality. The evaluation of relevance can be conducted using the SentenceTransformers framework (Reimers and Gurevych, 2019). Additionally, to evaluate hints for their readability, factors such as hint length and the difficulty of words can be considered as well as existing readability indexes or machine learning solutions could be harnessed (Martinc et al., 2021). Finally, assessing the ambiguity attribute proves challenging, and finding a suitable evaluation method is inherently difficult. Relevance and readability will be the focus of our future work.
@@ 4.1. Convergence Quality Attribute
Figure 4. Convergence Evaluator: The process begins by directing the question to the Candidate Generator stage, which generates up to twenty candidate answers. Subsequently, each generated candidate answer and the hint undergo an evaluation to determine the validity of the hint for the respective candidate answer. Finally, the results are conveyed to the Scoring stage for computation of the final score.
@@ 4.1.1. Candidate Generator
  Generate up to 20 candidate answers words for the question
  "QUESTION" in bullet points.
@@ 4.1.3. Scoring
In this module, we utilize the following equation to calculate the final score for the convergence quality attribute of the hint. To assess the convergence quality, if the Candidate Evaluator module identifies that the hint is not valid for the answer, we assign a score of zero, indicating that the hint is not useful for finding the answer. However, if the hint is valid for the answer, we compute the score based on the results for other candidate answers. A score of zero denotes that the hint does not reduce the number of candidate answers, whereas a score of one indicates that the hint successfully eliminates all candidate answers except for the correct one.
@@ 4.2. Familiarity Quality Attribute
We aim to measure the familiarity of information expressed in hints. Here, we use the popularity of the mentioned entities as a proxy for familiarity. The presence of popular entities can assist users in locating the answer more effortlessly, while less popular entities in a hint may potentially lead the user astray in finding the correct answer (Piao, 2021).
@@ 5.2. Human Evaluation
We engage human evaluators to assess the quality of the generated hints. To accomplish this, we randomly selected 100 questions from the test set, distributed as 25 questions from each coarse-grained class of question types (Entity, Human, Location, and Other), consisting of 9 labeled as hard, 8 labeled as medium, and 8 labeled as easy. In addition to the TriviaHG hints, we instruct LLaMA [^15] Please note that we employ LLaMA version 2 for this study. 7b (Touvron et al., 2023) vanilla to generate $10$ hints for each question [^16] We utilize the identical prompt as in Bing Chat AI for generating hints.. Furthermore, we finetune another LLaMA 7b instance on the training set of TrivisHG and then prompt it to generate $10$ hints for each question [^17] We perform model finetuning using the API functions available on AnyScale.com. Subsequently, we employ the hint filtering stage (Section 3.2.2) to refine the generated hints. At last, we obtain a total of 2,741 hints, comprising 950 hints from the TriviaHG dataset, 901 hints from the finetuned LLaMA 7b, and 890 hints from the LLaMA 7b vanilla. These data undergo evaluation in two distinct phases:
In the first phase, we ask $10$ human evaluators to evaluate the generated hints, providing ratings for attributes such as relevance, readability, ambiguity, convergence, and familiarity. Ratings are assigned on a scale of 1 to 5, where 1 denotes the lowest and 5 signifies the highest. They are also asked to search for each hint on both Google and Bing search engines, quickly examine the result page within a 5-second timeframe, and then determine and declare whether they could find the answer on the search results page. We request this action as it also can signal the quality of the hints. The reason is that it is plausible that a hint is still effective, while the knowledge of the annotators might be insufficient to answer the question utilizing the hint. The results of the human evaluation are summarized in Table 3, displaying average scores for each quality attribute. The relevance analysis reveals that hints generated by Bing Chat AI and the finetuned LLaMA 7b are relevant to the questions, while those from LLaMA 7b vanilla lack in terms of relevance. Additionally, the assessments of readability and ambiguity indicate that the generated hints are clear and comprehensible. The analysis of convergence and familiarity indicates that Bing Chat AI exhibits higher convergence and familiarity than other methods, while the finetuned LLaMA 7b demonstrates superior quality compared to the LLaMA 7b vanilla. Furthermore, Table 4 provides statistics on hints that annotators could use to find answers to the respective questions through searches on Google and Bing search engines.
@@
In the second phase, six distinct individuals, entirely separate from those involved in the first phase, are tasked with answering the questions using the provided hints. We randomly selected hints for each question from Bing Chat AI, LLaMA 7b vanilla, and finetuned LLaMA 7b. The procedure is as follows:
@@
Figure 8 provides an overview of the questions successfully answered by participants. The data indicates the high effectiveness of hints, as participants were able to answer only 19 questions before consulting hints, whereas after reviewing the hints, they could successfully answer 67 questions. The results from human evaluations underscore the superior efficacy of Bing Chat AI’s hints in aiding users in formulating answers compared to other methods. Our findings demonstrate that finetuned LLaMA 7b hints offer more value than LLaMA 7b vanilla hints, suggesting that leveraging the TriviaHG dataset for LLM finetuning contributes to the generation of superior hints. It is noteworthy that only $36\%$ of questions with hard answers were answerable, while medium and easy answers boasted success rates of $78\%$ and $96\%$, respectively. This observation emphasizes the increased challenge in generating hints for questions with difficult answers compared to their easier counterparts. Moreover, Figure 8 highlights that $64\%$ of the questions could be answered using Bing Chat AI hints, while the respective figures for the finetuned LLaMA 7b and LLaMA 7b vanilla are $48\%$ and $31\%$. It demonstrates the quality disparity between Bing Chat AI hints and the ones by the finetuned and vanilla LLaMA 7b.

### v2

@@ Abstract.
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 47th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 14–18, 2024; Washington, DC, USAbooktitle: Proceedings of the 47th International ACM SIGIR Conference on Research and Development in Information Retrieval (SIGIR ’24), July 14–18, 2024, Washington, DC, USAdoi: 10.1145/3626772.3657855isbn: 979-8-4007-0431-4/24/07ccs: Information systems Retrieval models and rankingccs: Information systems Evaluation of retrieval results
@@ 1. Introduction
Automatic question answering (QA) systems (Abdel-Nabi et al., 2023; Rogers et al., 2023; Mavi et al., 2022) have recently gained significant attention, allowing users to pose questions and receive direct answers. Traditional QA systems excel in specific knowledge domains by extracting relevant text spans from passages. In contrast, Large Language Models (LLMs) (Team et al., 2023; Workshop et al., 2022; Brown et al., 2020a), pre-trained on vast datasets, offer a more versatile approach. LLMs can handle multiple NLP tasks, including automatic QA, and can answer questions across diverse domains. Unlike extractive QA systems, they generate answers by creating new content, showcasing high proficiency and potential as an alternative to traditional QA systems.
Figure 1. Example hints for two sample questions in TriviaHG with their computed convergence quality (HICOS) and familiarity (HIFAS) provided on a scale of 1 to 5 (1 is the lowest and 5 is the highest quality).
@@
The use of hints to assist users is not a new concept in computer science. For example, in the field of automatic hint generation for programming (AHGP)  (McBroom et al., 2021), the task is to generate hints automatically to aid students in coding exercises. However, there is limited research on automatic hint generation specifically for factoid questions (AHGQ) - questions that can be answered with facts expressed in a few words (Jatowt et al., 2023). While both AHGP and AHGQ aim to assist humans, the former is more tailored to coding exercises, often suggesting code edits, and the latter is a more general task applicable across various types of questions. Furthermore, there is an additional benefit from focusing on AHGQ task. Hints for questions could be used to enhance Retrieval Augmented Generation (RAG) solutions (Gao et al., 2023; Abdallah and Jatowt, 2023) and Query Expansion (Wang et al., 2023) methods. Specifically, in the context of Retrieval Augmented Generation, incorporating hints as query explanations has the potential to enhance the model’s ability to understand and process queries.As mentioned above, the hints may also find applications in entertainment, such as in game shows like Jeopardy.
As a novel task, AHGQ faces challenges such as the absence of datasets and reliable evaluation methods. Addressing this gap, we introduce a framework for AHGQ and utilize it to generate TriviaHG [^1] The code, dataset, and experimental results are freely available at https://github.com/DataScienceUIBK/TriviaHG, a dataset for automatic hint generation for factoid questions. Our approach relies on web pages as the primary information source and utilizes Copilot (formerly known as Bing Chat AI) [^2] https://copilot.microsoft.com to generate hints based on online sources, ensuring accuracy and relevance. The hints are designed to be user-friendly and helpful.
As a second contribution, we propose an automatic evaluation method to assess hint quality, considering aspects of convergence and familiarity. The suggested evaluation method is versatile, suitable for application across various types of hints, regardless of their format, whether they be in documents, keywords, or other forms. The convergence quality is a characteristic that reflects the extent to which a hint can narrow down or eliminate potential answers to a given question. The familiarity is a characteristic that signifies the degree to which the entities mentioned in a hint are well-known or widely recognized. We label the score of convergence quality as HICOS (HInt COnvergence Score), and the score of familiarity quality as HIFAS (HInt FAmiliarity Score). Figure 1 illustrates two questions along with a few examples hints from TriviaHG along with the computed HICOS and HIFAS.
@@ 2.1. QA Task
The prior efforts have not focused on generating hints but rather on providing the final answers.
@@ 2.3. Evaluation Methods
Evaluation methods for AHGP can be broadly classified into four categories: (1) User Studies and Surveys, (2) Comparisons with Experts, (3) Analysis Using Historical Data, and (4) Technical Evaluations (McBroom et al., 2021). The User Studies and Surveys category involves collecting feedback from users (typically students) who interact with the hinting system (Head et al., 2017; Lee et al., 2018). The Comparisons with Experts category involves comparing the hints generated by the system with those provided by human experts (Piech et al., 2015; Price et al., 2017). The Analysis Using Historical Data involves using past student interaction data to evaluate the effectiveness of hints (Lazar et al., 2017; Haldeman et al., 2018). The Technical Evaluations category involves assessing the system’s performance in terms of computational efficiency, scalability, and robustness (Marin et al., 2017).
All the evaluation methods for QA are automatic, but there are a few automatic evaluation methods for AHGP. It shows that assessing hint quality automatically is a complex task. However, the automatic evaluation methods for AGHP cannot be employed to assess AGHQ because they just focus on the programming aspects of hints. As said, Jatowt et al. (2023) is the only study focused on the AGHQ, but they also evaluate the quality of hints manually by asking humans to assess the hints and do not propose any automatic evaluation methods for AGHQ. In our work, we introduce a fully automated method for evaluating hint qualities.
@@ 3.1.1. Question Selection
We select the TriviaQA dataset (Joshi et al., 2017) as our primary question source because it offers challenging trivia-style questions covering a wide range of diverse topics. It has also been frequently used in QA research (Beltagy et al., 2020; Zaheer et al., 2020; Su et al., 2024; Prasad et al., 2023; Nassiri and Akhloufi, 2023). We filter out questions with fewer than 6 words or exceeding 20 words, as well as those lacking question marks. Additionally, questions whose answers do not have corresponding Wikipedia pages are excluded, as Wikipedia is utilized for evaluating hint familiarity quality. After applying these criteria, 66,300 questions remain available for further processing.
@@ 3.2.1. Hint Generator
We utilize LLMs to generate hints for the questions, specifically selecting Copilot as our preferred choice. Copilot functions as a RAG system, leveraging also information from retrieved documents on the Internet to formulate answers to user questions. This system does not only generate answers based on existing online documents but also provides the sources of the hints and answers. In contrast, LLMs like GPT-4 (Achiam et al., 2023), Gemini (Team et al., 2023), and WizardLM (Xu et al., 2023) rely on their stored knowledge to generate answers without indicating the sources. Therefore, Copilot is considered more reliable for hint generation as it transparently presents the sources of the generated hints. To generate hints for questions, we follow these steps:
@@
    We first prompt Copilot to answer a question. This enables us to determine if Copilot is capable of retrieving the answer. If it is unable to do so, we discard the question.
@@
    We instruct Copilot to produce 10 hints for the question without including the answer in the hints. This approach allows us to generate hints without the risk of answer leakage, i.e., the case when a hint explicitly reveals the answer to the question in its content [^6] For instance, for the question Who won the FIVB Volleyball Men’s World Championship in 2018?, the hint Their official language is Polish. exhibits answer leakage.. Figure 3 illustrates the trend of using prompts in Copilot for hint generation.
Figure 3. The hint generation system initiates by prompting the question to Copilot. Then, it produces a snippet serving as the answer. Following this, we assess the correctness of the provided answer. If the answer is correct, we prompt Copilot to generate 10 hints. The numbers in the brackets denote the source pages for both the answer and the hints.
@@ 3.2.2. Hint Filtering
While we advise Copilot to exercise caution regarding the problem of answer leakage, there are instances where it might produce hints inadvertently revealing the answer. Additionally, there are occasions when it generates hints that merely rephrase or describe the questions. For instance, for the question In which city are the headquarters of the International Monetary Fund?, the model may generate a hint like The headquarters of the International Monetary Fund is located in this city. Such hints do not help users in finding correct answers. To address these issues, we take the following steps:
@@
Convergence The value of HICOS
Familiarity The value of HIFAS
3.2.3. Dataset Attributes
@@ 4. Automatic Evaluation Method
In this work, we focus on automating the evaluation of convergence and familiarity quality attributes, as, in our opinion, they are the strongest indicators of hint quality. The evaluation of relevance can be conducted using, for example, the SentenceTransformers framework (Reimers and Gurevych, 2019). Additionally, to evaluate hints for their readability, factors such as hint length and the difficulty of words can be considered as well as existing readability indexes or machine learning solutions could be harnessed (Martinc et al., 2021). Finally, assessing the ambiguity attribute proves challenging, and finding a suitable evaluation method may be difficult. Relevance and readability will be the focus of our future work.
@@ 4.1. Convergence Quality Attribute
Figure 4. Convergence Evaluator: The process begins by directing the question to the Candidate Generator stage, which generates up to twenty candidate answers. Subsequently, each generated candidate answer and the hint undergo an evaluation to determine the validity of the hint for the respective candidate answer. Finally, the results are conveyed to the Scoring stage for computation of the HICOS.
@@ 4.1.1. Candidate Generator
  Generate up to 20 candidate answers words for the quest
  ion "QUESTION" in bullet points.
@@ 4.1.3. Scoring
In this module, we utilize the following equation to calculate the HICOS of the hint. To assess the convergence quality, if the Candidate Evaluator module identifies that the hint is not valid for the answer, we assign a score of zero, indicating that the hint is not useful for finding the answer. However, if the hint is valid for the answer, we compute the score based on the results for other candidate answers. A score of zero denotes that the hint does not reduce the number of candidate answers, whereas a score of one indicates that the hint successfully eliminates all candidate answers except for the correct one.
@@ 4.2. Familiarity Quality Attribute
We aim to measure the familiarity of information expressed in hints (HIFAS). Here, we use the popularity of the mentioned entities as a proxy for familiarity. The presence of popular entities can assist users in locating the answer more effortlessly, while less popular entities in a hint may potentially lead the user astray in finding the correct answer (Piao, 2021).
@@ 5.2. Human Evaluation
We engage human evaluators to assess the quality of the generated hints. To accomplish this, we randomly selected 100 questions from the test set, distributed as 25 questions from each coarse-grained class of question types (Entity, Human, Location, and Other), consisting of 9 labeled as hard, 8 labeled as medium, and 8 labeled as easy. In addition to the TriviaHG hints, we instruct LLaMA [^15] Please note that we employ LLaMA version 2 for this study. 7b (Touvron et al., 2023) vanilla to generate $10$ hints for each question [^16] We utilize the identical prompt as in Copilot for generating hints.. Furthermore, we finetune another LLaMA 7b instance on the training set of TriviaHG and then prompt it to generate $10$ hints for each question [^17] We perform model finetuning using the API functions available on AnyScale.com. Subsequently, we employ the hint filtering stage (Section 3.2.2) to refine the generated hints. At last, we obtain a total of 2,741 hints, comprising 950 hints from the TriviaHG dataset, 901 hints from the finetuned LLaMA 7b, and 890 hints from the LLaMA 7b vanilla. These data undergo evaluation in two distinct phases:
In the first phase, we ask $10$ human evaluators to evaluate the generated hints, providing ratings for attributes such as relevance, readability, ambiguity, convergence, and familiarity. Ratings are assigned on a scale of 1 to 5, where 1 denotes the lowest and 5 signifies the highest. They are also asked to search for each hint on both Google and Bing search engines, quickly examine the result page within a 5-second timeframe, and then determine and declare whether they could find the answer on the search results page. We request this action as it also can signal the quality of the hints. The reason is that it is plausible that a hint is still effective, while the knowledge of the annotators might be insufficient to answer the question utilizing the hint. The results of the human evaluation are summarized in Table 3, displaying average scores for each quality attribute. The relevance analysis reveals that hints generated by Copilot and the finetuned LLaMA 7b are relevant to the questions, while those from LLaMA 7b vanilla lack in terms of relevance. Additionally, the assessments of readability and ambiguity indicate that the generated hints are clear and comprehensible. The analysis of convergence and familiarity indicates that Copilot exhibits higher convergence and familiarity than other methods, while the finetuned LLaMA 7b demonstrates superior quality compared to the LLaMA 7b vanilla. Furthermore, Table 4 provides statistics on hints that annotators could use to find answers to the respective questions through searches on Google and Bing search engines.
@@
In the second phase, six distinct individuals, entirely separate from those involved in the first phase, are tasked with answering the questions using the provided hints in a similar way to the approach in (Jatowt et al., 2023). We randomly select hints for each question from Copilot, LLaMA 7b vanilla, and finetuned LLaMA 7b. The procedure is as follows:
@@
Figure 8 provides an overview of the questions successfully answered by participants. The data indicates the high effectiveness of hints, as participants could answer only 19 questions before consulting hints, whereas after reviewing the hints, they could successfully answer 67 questions. The results from human evaluations underscore the superior efficacy of Copilot’s hints in aiding users in formulating answers compared to other methods. Our findings show that finetuned LLaMA 7b hints offer more value than LLaMA 7b vanilla hints, suggesting that leveraging the TriviaHG dataset for LLM finetuning contributes to the generation of superior hints. Only $36\%$ of questions with hard answers were answerable, while medium and easy answers boasted success rates of $78\%$ and $96\%$, respectively. This observation emphasizes the increased challenge in generating hints for questions with difficult answers compared to their easier counterparts. Moreover, Figure 8 highlights that $64\%$ of the questions could be answered using Copilot hints, while the respective figures for the finetuned LLaMA 7b and LLaMA 7b vanilla are $48\%$ and $31\%$. It demonstrates the quality disparity between Copilot and the finetuned and vanilla LLaMA 7b hints.